# Learning Useful Representations of Recurrent Neural Network Weight Matrices

### v1

@@ Abstract
Recurrent Neural Networks (RNNs) are general-purpose parallel-sequential computers. The program of an RNN is its weight matrix. How to learn useful representations of RNN weights that facilitate RNN analysis as well as downstream tasks? While the mechanistic approach directly looks at some RNN’s weights to predict its behavior, the functionalist approach analyzes its overall functionality—specifically, its input-output mapping. We consider several mechanistic approaches for RNN weights and adapt the permutation equivariant Deep Weight Space layer for RNNs. Our two novel functionalist approaches extract information from RNN weights by ‘interrogating’ the RNN through probing inputs. We develop a theoretical framework that demonstrates conditions under which the functionalist approach can generate rich representations that help determine RNN behavior. We create and release the first two ‘model zoo’ datasets for RNN weight representation learning. One consists of generative models of a class of formal languages, and the other one of classifiers of sequentially processed MNIST digits. With the help of an emulation-based self-supervised learning technique we compare and evaluate the different RNN weight encoding techniques on multiple downstream applications. On the most challenging one, namely predicting which exact task the RNN was trained on, functionalist approaches show clear superiority.
@@ 3 RNN Weight Encoders
When using NN weights as input for another network, two main challenges arise. First, the number of weights can quickly become very large. Second, the weight space exhibits symmetries, particularly with respect to the permutation of hidden neurons (Hecht-Nielsen, 1990). Rearranging the order of hidden neurons in a network does not change the computation it performs [^2] With piece-wise linear activation functions like ReLU, there is also invariance to certain types of weight scaling. However, this is not a focus here as RNNs usually use different activation functions.. An effective NN weight encoder should recognize these symmetries.
We consider an RNN, $f_{\theta}:\mathbb{R}^{X}\times\mathbb{R}^{H}\to\mathbb{R}^{Y}\times\mathbb{R}^{H};(x,h_{o})\mapsto(y,h_{n})$, parametrized by $\theta\in\Theta$, which maps an input $x$ and hidden state $h_{o}$ to an output $y$ and a new hidden state $h_{n}$. In the following, we assume that all RNNs are multi-layer LSTMs [^3] Generalizing our framework to other RNN architectures should be straightforward.. An RNN weight encoder is a function $E_{\phi}:\Theta\to\mathbb{R}^{Z};\theta\mapsto z$, mapping RNN weights $\theta$ to $Z$-dimensional representation vectors $z=E_{\phi}(\theta)$. The encoder’s parameters are $\phi\in\Phi$.
@@
This approach, successfully used in previous studies (Unterthiner et al., 2020) to predict properties of CNNs, involves creating, for each weight matrix, a vector consisting of mean, standard deviations, and five quantiles ($0,0.25,0.5,0.75,1$). For LSTMs, each layer yields three distinct vectors corresponding to the input-to-hidden weights, hidden-to-hidden weights, and the bias vector (assuming that the weight matrices of the four LSTM gates are consolidated into one). These vectors are then concatenated and given to a multi-layer perceptron (MLP). The architecture is inherently invariant to permutations of hidden neurons. It efficiently scales with RNN size due to its reliance on high-level features. However, this invariance extends to many transformations beyond those that preserve the RNN functionality. Consequently, RNNs with identical layer-wise statistics may behave differently, preventing the encoder’s ability to approximate all functions of the weight space.
@@
In this approach, all RNN weights are flattened into a single vector before being fed into an MLP. Unlike the previous architecture, this one is not invariant to hidden neuron permutations. As a result, it faces a different challenge: numerous RNNs might appear different to the encoder yet perform identical computations. This impedes generalization, as the MLP has difficulty learning these symmetries (empirically demonstrated in Appendix F.2). Another issue is the very large size of the input vector. Let $N$ represent the number of hidden neurons in $f_{\theta}$. The number of parameters in the input layer of the MLP is proportional to $N^{2}$. However, with an adequately large MLP, this architecture can approximate any function in the weight space [^4] When mentioning universality of weight space functions $\Theta\to\mathbb{R}^{Z}$, we imply the regularity conditions of Navon et al. (2023), Proposition 6.2.. This follows immediately from the universal approximation property of MLPs (Hornik, 1991).
@@
Both Navon et al. (2023) and Zhou et al. (2023) proposed architectures for processing the weights of feedforward NNs, closely related in design. These architectures are invariant precisely to permutations of hidden neurons and are capable of universally approximating functions of the weight space. We refer to this architecture as DWSNet, following Navon et al. (2023), and extend its application to LSTM networks. The central concept of DWSNet is to construct layers that are equivariant to the hidden neuron permutation group. These layers process the weights, where each weight is represented by a feature vector. A final pooling layer across all weights ensures invariance to neuron permutation. The mechanisms for adapting DWSNets to LSTMs, along with arguments for their universality, are presented in Appendix C. Appendix F.2 validates the implementation by demonstrating that DWSNets maintain invariance only to correctly permuted weights in LSTMs.
@@
The probing sequence length, denoted as $L$, is a fixed hyperparameter. A non-interactive probing encoder uses a sequence of learnable embeddings, $S_{i},1\leq i\leq L$. At each probing step $i$, the embedding $S_{i}$ is transformed by $E_{I}$, an MLP followed by a reshaping operation that outputs a set of $M$ probing inputs $\hat{x}_{im}\in\mathbb{R}^{X}\;\forall m\in\{1,\dots,M\}$. $M$ represents the number of parallel probing sequences. These inputs are processed as a batch by the RNN $f_{\theta}$, yielding in a batch of probing outputs $\hat{y}_{im}=f_{\theta}(\hat{x}_{im})\in\mathbb{R}^{Y}$  [^5] For $f_{\theta}$, the recurrence of the hidden states over sequence steps is implied.. The probing outputs are concatenated and further transformed by another MLP $E_{O}$, resulting in a vector $o_{i}$. The sequence of vectors $o_{i}$ is processed by an LSTM network $E_{R}$, which then outputs the RNN representation $E(\theta)$.
@@
Both types of probing encoder retain the invariance properties of $f_{\theta}$. However, functionalist encoders have limitations in differentiating between weight space functions; they cannot discern mechanistic differences in functionally equivalent RNNs. Consequently, two RNNs performing exactly the same function will look identical to a probing encoder, even if they use different algorithms to compute the function. This means probing encoders are not universal in the sense that some of the mechanistic approaches are. Table 1 summarizes key properties of the different encoder architectures.
Table 1: Properties of the different RNN weight encoder architectures. $N$ is the number of hidden neurons in $f_{\theta}$. Encoder Permutation Invariant Universal Approx. #Params Type Layerwise Statistics Yes No const. Mechanistic Flattened Weights No Yes $O(N^{2})$ Mechanistic Parameter Transformer No Yes $O(N)$ Mechanistic DWSNet Yes Yes const. Mechanistic Non-Interactive Probing Yes No const. Functionalist Interactive Probing Yes No const. Functionalist
[Uncaptioned image] Figure 3: Interrogator $I_{D}$ has access to a set $D$ of functions and interacts with function $f_{C}$, which it has to identify.
@@ 3.1 Theoretical Aspects of the Functionalist Approach
We introduce a theoretical framework for analyzing probing encoders and the distinctions between interactive and non-interactive settings. In practice, any RNN weight encoder is trained on a finite number of distinct RNNs. For each of them, it should output a unique representation. In this section we study the related task of uniquely identifying a function from a given set by interacting with it. Rather than using RNNs, we examine total computable functions, which are functions that halt and produce an output for every input. This is a minor limitation, since in practical scenarios, RNNs are almost always given a finite runtime.
Let $D$ represent a set of $n$ total computable functions $\{f_{i}:\mathbb{N}\to\mathbb{N}|i=1,2,\ldots,n\}$. In other words, $D$ comprises $n$ Turing machines that halt on every input, with no pair being functionally equivalent. Let $I_{D}$ denote another Turing machine, which we call the Interrogator. $I_{D}$ has access to the function set $D$ (e.g., the corresponding Turing numbers might be written somewhere on its tape). Moreover, $I_{D}$ is given access to one function $f_{C}\in D$ as a black box. $I_{D}$ can interact with $f_{C}$ by providing an input $x\in\mathbb{N}$ and subsequently reading the corresponding output $f_{C}(x)$. The task of $I_{D}$ is to identify which member of $D$ corresponds to function $f_{C}$, while minimizing interactions with $f_{C}$. Specifically, $I_{D}$ must return $i\in\{1,\dots,n\}$ such that $f_{C}=f_{i}$. This setup is depicted in Figure 3. The proofs of the following propositions can be found in Appendix B.
@@ 4 Self-Supervised Learning of RNN Weight Representations
with fixed initial states $y_{0},\eta_{0}$ and $h_{0}$. For instance, $f_{\theta}$ might be an autoregressive generative model, with $\mathcal{E}$ acting as a stochastic environment that receives a probability distribution over some language tokens, $y_{t}$—the output of $f$—, and produces a representation (e.g., a one-hot vector) of the new input token $x_{t+1}$. When the environment is stochastic, numerous rollouts can be generated for any $\theta\in\Theta$. A rollout sequence of a function $f_{\theta}$ in environment $\mathcal{E}$ has the form $S_{\theta}=(x_{1},y_{1},x_{2},y_{2},\dots)$.
To train an RNN weight encoder $E_{\phi}$, we consider an Emulator $A_{\xi}:\mathbb{R}^{X}\times\mathbb{R}^{B}\times\mathbb{R}^{Z}\to\mathbb{R}^{Y}\times\mathbb{R}^{B};(x,b_{o},z)\mapsto(\tilde{y},b_{n})$, parametrized by $\xi\in\Xi$. The Emulator is an RNN with hidden state $b$ that learns to imitate different RNNs $f_{\theta}$ based on their function encoding $z=E(\theta)$.
We consider a dataset $\mathcal{D}=\{(\theta_{i},S_{\theta_{i}})|i=1,2,\dots\}$ composed of tuples, each containing the parameters of a different RNN and a corresponding rollout sequence. We assume that all RNNs have the same initial state $h_{0}$ but have been trained on different tasks. The Encoder $E_{\phi}$ and the Emulator $A_{\xi}$ are jointly trained by minimizing a loss function $\mathcal{L}$. This loss function measures the behavioral similarity between an RNN $f_{\theta}$ and the Emulator $A_{\xi}$, which is conditioned on the function representation $z=E_{\phi}(\theta)$ of $\theta$ as produced by the Encoder $E_{\phi}$ (see Figure 4). Put simply, the Emulator uses the representations of a set of diverse RNNs $f_{\theta}$ to imitate their behavior: [^6] The recurrence of $A_{\xi}$ is omitted for clarity.
@@ 5.1 Formal Languages
All models are trained on sequences of length 42, including one begin-of-sequence (BOS) and one end-of-sequence (EOS) token. The maximum value of $n$ is 10. If a language string is shorter than 42, it is padded at the end with EOS tokens. The OOD test set comprises RNNs trained on languages where the sum of the absolute values of $m_{b},m_{c}$ and $m_{d}$ is the smallest.
@@ 6 Experiments and Results
Our empirical investigation comprises two experimental phases. In the first phase, we apply the emulation-based representation method described in Section 6.1 to learn representations for RNNs from formal languages and the sequential MNIST dataset. The second phase is dedicated to predicting properties of the RNNs. These predictions are either based on representations learned in the first phase or derived from fully supervised models trained from scratch. We conduct the main experiments using 15 different random seeds for each model. The outcomes are presented as bootstrapped means with 95% confidence intervals. For RNNs trained on the Formal Languages dataset, performance is measured by the proportion of correctly generated strings (i.e., strings belonging to the language on which it was trained). For the Sequential MNIST dataset, we assess performance using standard digit classification validation accuracy. For the Flattened Weights and the Parameter Transformer encoders, the training data is augmented by randomly permuting the neurons of the input RNN.
@@ 6.1 Representation Learning
Layer-Wise Statistics 0.048 $\scriptstyle{(0.048,0.049)}$ 0.041 $\scriptstyle{(0.040,0.042)}$
@@
Figure 5: $f_{\theta}$’s original performance on formal language generation vs. the performance of $A_{\xi}$’s emulation based on $E_{\phi}(\theta)$.
The six different encoder architectures are trained according to Objective 1. The hyperparameters of these encoders are selected to ensure a comparable number of parameters across all models. Each encoder generates a 16-dimensional representation $z$. An LSTM with two layers functions as the emulator $A_{\xi}$. The conditioning of $A_{\xi}$ on an RNN $f_{\theta}$ is implemented by incorporating a linear projection of the corresponding representation $z$ to the BOS token of the input sequence of $A_{\xi}$. More details and hyperparameters can be found in Appendix E.
Table 2 displays the emulation losses for the different encoder types on both the formal languages and the sequential MNIST validation datasets. For formal languages, the interactive probing method outperforms the others. In the case of sequential MNIST, the performance of both probing encoders is quite similar and superior to that of the mechanistic encoders. Figure 5 shows the emulation effectiveness of various RNNs from the formal languages dataset. Each marker represents an RNN $f_{\theta}$. The x-axis indicates $f_{\theta}$’s performance on its trained language, while the y-axis reflects $A_{\xi}$’s performance on the same language when conditioned on $E_{\phi}(\theta)$. In this dataset, only the interactive probing encoder yields representations that enable $A_{\xi}$ to effectively emulate the original RNN. For the sequential MNIST dataset, however, emulation performance is more uniform across encoder types, as depicted in Figure 14 in the Appendix. Figure 7 examines the structure of the embedding space created by the interactive probing encoder for the formal languages validation set. Each point represents an RNN, with distinct clustering observed for each specific formal language. We highlight one cluster, corresponding to the language $L_{-2,2,-2}$. Each cluster shows a gradient representing the generation accuracy of the RNNs. Figure 7 shows the embedding space for the sequential MNIST dataset, where the task involves rotating MNIST digits. This rotation, a continuous scalar, is evident in the embedding space, alongside the classification accuracy of each model. For contrast, a t-SNE dimensionality reduction applied directly to the model weights is shown. Although small clusters are noticeable for the nine snapshots of each run, the overall embedding space lacks a coherent structure. Visualizations of embedding spaces for all encoder architectures and OOD test sets are presented in Appendix F.4.

### v2

@@ Abstract
Recurrent Neural Networks (RNNs) are general-purpose parallel-sequential computers. The program of an RNN is its weight matrix. How to learn useful representations of RNN weights that facilitate RNN analysis as well as downstream tasks? While the mechanistic approach directly looks at some RNN’s weights to predict its behavior, the functionalist approach analyzes its overall functionality—specifically, its input-output mapping. We consider several mechanistic approaches for RNN weights and adapt the permutation equivariant Deep Weight Space layer for RNNs. Our two novel functionalist approaches extract information from RNN weights by ‘interrogating’ the RNN through probing inputs. We develop a theoretical framework that demonstrates conditions under which the functionalist approach can generate rich representations that help determine RNN behavior. We release the first two ‘model zoo’ datasets for RNN weight representation learning. One consists of generative models of a class of formal languages, and the other one of classifiers of sequentially processed MNIST digits. With the help of an emulation-based self-supervised learning technique we compare and evaluate the different RNN weight encoding techniques on multiple downstream applications. On the most challenging one, namely predicting which exact task the RNN was trained on, functionalist approaches show clear superiority.
@@ 3 RNN Weight Encoders
When using NN weights as input for another network, two main challenges arise. First, the number of weights can quickly become very large. Second, the weight space exhibits symmetries, particularly with respect to the permutation of hidden neurons (Hecht-Nielsen, 1990). Rearranging the order of hidden neurons in a network does not change the computation it performs [^2] With piece-wise linear activation functions like ReLU, there is also invariance to certain types of weight scaling. However, this is not a focus here as RNNs usually use different activation functions. There is another symmetry with respect to sign flips.. An effective NN weight encoder should recognize these symmetries.
We consider an RNN, $f_{\theta}:\mathbb{R}^{X}\times\mathbb{R}^{H}\to\mathbb{R}^{Y}\times\mathbb{R}^{H};(x,h_{t-1})\mapsto(y,h_{t})$, $t\in 1,2,\dots$, parametrized by $\theta\in\Theta$, which maps an input $x$ and hidden state $h_{t-1}$ to an output $y$ and a new hidden state $h_{t}$. In the following, we assume that all RNNs are multi-layer LSTMs [^3] Generalizing our framework to other RNN architectures should be straightforward.. An RNN weight encoder is a function $E_{\phi}:\Theta\to\mathbb{R}^{Z};\theta\mapsto z$, mapping RNN weights $\theta$ to $Z$-dimensional representation vectors $z=E_{\phi}(\theta)$. The encoder’s parameters are $\phi\in\Phi$.
@@
This approach, successfully used in previous studies (Unterthiner et al., 2020) to predict properties of CNNs, involves creating, for each weight matrix, a vector consisting of mean, standard deviations, and five quantiles ($0,0.25,0.5,0.75,1$). For LSTMs, each layer yields twelve distinct vectors, for each of the four gates corresponding to the input-to-hidden weights, hidden-to-hidden weights, and the bias vector. These vectors are then concatenated and given to a multi-layer perceptron (MLP). The architecture is inherently invariant to permutations of hidden neurons. It efficiently scales with RNN size due to its reliance on high-level features. However, this invariance extends to many transformations beyond those that preserve the RNN functionality. Consequently, RNNs with identical layer-wise statistics may behave differently, preventing the encoder’s ability to approximate all functions of the weight space.
@@
In this approach, all RNN weights are flattened into a single vector before being fed into an MLP. Unlike the previous architecture, this one is not invariant to hidden neuron permutations. As a result, it faces a different challenge: numerous RNNs might appear different to the encoder yet perform identical computations. This impedes generalization, as the MLP has difficulty learning these symmetries (empirically demonstrated in Appendix E.2). Another issue is the very large size of the input vector. Let $N$ represent the number of hidden neurons in $f_{\theta}$. The number of parameters in the input layer of the MLP is proportional to $N^{2}$. However, with an adequately large MLP, this architecture can approximate any function in the weight space [^4] When mentioning universality of weight space functions $\Theta\to\mathbb{R}^{Z}$, we imply the regularity conditions of Navon et al. (2023), Proposition 6.2.. This follows immediately from the universal approximation property of MLPs (Hornik, 1991).
@@
Both Navon et al. (2023) and Zhou et al. (2023) proposed architectures for processing the weights of feedforward NNs, closely related in design. These architectures are invariant precisely to permutations of hidden neurons and are capable of universally approximating functions of the weight space. We refer to this architecture as DWSNet, following Navon et al. (2023), and extend its application to LSTM networks. The central concept of DWSNet is to construct layers that are equivariant to the hidden neuron permutation group. These layers process the weights, where each weight is represented by a feature vector. A final pooling layer across all weights ensures invariance to neuron permutation. The mechanisms for adapting DWSNets to LSTMs, along with arguments for their universality, are presented in Appendix B. Appendix E.2 validates the implementation by demonstrating that DWSNets maintain invariance only to correctly permuted weights in LSTMs.
@@
The probing sequence length, denoted as $L$, is a fixed hyperparameter. A non-interactive probing encoder uses a sequence of learnable embeddings, $S_{i},1\leq i\leq L$. At each probing step $i$, the embedding $S_{i}$ is transformed by $E_{I}$, an MLP followed by a reshaping operation that outputs a set of $M$ probing inputs $\hat{x}_{im}\in\mathbb{R}^{X}\;\forall m\in\{1,\dots,M\}$. $M$ represents the number of parallel probing sequences. These inputs are processed as a batch by the RNN $f_{\theta}$, resulting in a batch of probing outputs $\hat{y}_{im}=f_{\theta}(\hat{x}_{im})\in\mathbb{R}^{Y}$  [^5] For $f_{\theta}$, the recurrence of the hidden states over sequence steps is implied.. The probing outputs are concatenated and further transformed by another MLP $E_{O}$, resulting in a vector $o_{i}$. The sequence of vectors $o_{i}$ is processed by an LSTM network $E_{R}$, which then outputs the RNN representation $E(\theta)$.
@@
Both types of probing encoders retain the invariance properties of $f_{\theta}$. However, functionalist encoders have limitations in differentiating between weight space functions; they cannot discern mechanistic differences in functionally equivalent RNNs. Consequently, two RNNs performing exactly the same function will look identical to a probing encoder, even if they use different algorithms to compute the function. This means probing encoders are not universal in the sense that some of the mechanistic approaches are. Table 1 summarizes key properties of the different encoder architectures.
Table 1: Properties of the different RNN weight encoder architectures. $N$ is the number of hidden neurons in $f_{\theta}$.
Encoder
Permutation
Invariant
Universal
Approx.
#Params Type
Layerwise Statistics Yes No const. Mechanistic
Flattened Weights No Yes $O(N^{2})$ Mechanistic
Parameter Transformer No Yes $O(N)$ Mechanistic
DWSNet Yes Yes const. Mechanistic
Non-Interactive Probing Yes No const. Functionalist
Interactive Probing Yes No const. Functionalist
Figure 3: Interrogator $I_{D}$ has access to a set $D$ of functions and interacts with function $f_{C}$, which it has to identify.
@@ 3.1 Theoretical Aspects of the Functionalist Approach
We introduce a theoretical framework for analyzing probing encoders and the distinctions between interactive and non-interactive settings. In practice, any RNN weight encoder is trained on a finite number of distinct RNNs. For each of them, it should output a unique representation. In this section, we study the related task of uniquely identifying a function from a given set by interacting with it. Rather than using RNNs, we examine total computable functions, which are functions that halt and produce an output for every input. This is a minor limitation, since in practical scenarios, RNNs are almost always given a finite runtime.
Let $D$ represent a set of $n$ total computable functions $\{f_{i}:\mathbb{N}\to\mathbb{N}|i=1,2,\ldots,n\}$. In other words, $D$ comprises $n$ Turing machines that halt on every input, with no pair being functionally equivalent. Let $I_{D}$ denote another Turing machine, which we call the Interrogator. $I_{D}$ has access to the function set $D$ (e.g., the corresponding Turing numbers might be written somewhere on its tape). Moreover, $I_{D}$ is given access to one function $f_{C}\in D$ as a black box. $I_{D}$ can interact with $f_{C}$ by providing an input $x\in\mathbb{N}$ and subsequently reading the corresponding output $f_{C}(x)$. The task of $I_{D}$ is to identify which member of $D$ corresponds to function $f_{C}$, while minimizing interactions with $f_{C}$. Specifically, $I_{D}$ must return $i\in\{1,\dots,n\}$ such that $f_{C}=f_{i}$. This setup is depicted in Figure 3. It should be mentioned that RNNs of finite size and precision are not universal computers (Merrill, 2019; Delétang et al., 2022). However, the following propositions depend mainly on the relative computational ability of $I_{D}$ and the functions in $D$. Hence we choose this simple abstract framework of distinguishing total computable functions and believe that it transfers well into realistic RNN settings. The proofs of the following propositions can be found in Appendix A.
@@ 4 Self-Supervised Learning of RNN Weight Representations
with fixed initial states $y_{0},\eta_{0}$ and $h_{0}$. For instance, $f_{\theta}$ might be an autoregressive generative model, with $\mathcal{E}$ acting as a stochastic environment that receives a probability distribution over some language tokens, $y_{t}$—the output of $f$ at time step $t$—, and produces a representation (e.g., a one-hot vector) of the new input token $x_{t+1}$. When the environment is stochastic, numerous rollouts can be generated for any $\theta\in\Theta$. A rollout sequence of a function $f_{\theta}$ in environment $\mathcal{E}$ has the form $S_{\theta}=(x_{1},y_{1},x_{2},y_{2},\dots)$.
To train an RNN weight encoder $E_{\phi}$, we consider an Emulator $A_{\xi}:\mathbb{R}^{X}\times\mathbb{R}^{B}\times\mathbb{R}^{Z}\to\mathbb{R}^{Y}\times\mathbb{R}^{B};(x,b_{t-1},z)\mapsto(\tilde{y},b_{t})$, parametrized by $\xi\in\Xi$. The Emulator is an RNN with hidden state $b$ that learns to imitate different RNNs $f_{\theta}$ based on their function encoding $z=E(\theta)$.
We consider a dataset $\mathcal{D}=\{(\theta_{i},S_{\theta_{i}})|i=1,2,\dots\}$ composed of tuples, each containing the parameters of a different RNN and a corresponding rollout sequence. We assume that all RNNs have the same initial state $h_{0}$ but have been trained on different tasks. The Encoder $E_{\phi}$ and the Emulator $A_{\xi}$ are jointly trained by minimizing a loss function $\mathcal{L}$. This loss function measures the behavioral similarity between an RNN $f_{\theta}$ and the Emulator $A_{\xi}$, which is conditioned on the function representation $z=E_{\phi}(\theta)$ of $\theta$ as produced by the Encoder $E_{\phi}$ (see Figure 4). Put simply, the Emulator uses the representations of a set of diverse RNNs $f_{\theta}$ to imitate their behavior: [^6] The recurrence of $A_{\xi}$ is omitted for simplicity.
@@
If we are interested only in the RNN weight encoder $E_{\phi}$, the emulator $A_{\xi}$ can be discarded after training. However, there might be potential applications for the emulator, for example, in the context of imitation learning (Zare et al., 2023), behavior cloning (Torabi et al., 2018), or consolidating the knowledge from many different models (Joseph & N Balasubramanian, 2020).
@@ 5.1 Formal Languages
All models are trained on sequences of length 42, including one begin-of-sequence (BOS) and one end-of-sequence (EOS) token. The maximum value of $n$ is 10. If a language string is shorter than 42, it is padded at the end with EOS tokens. The OOD test set contains the RNNs trained on languages where the sum of the absolute values of $m_{b},m_{c}$ and $m_{d}$ is the smallest.
@@ 6 Experiments and Results
Our empirical investigation involves two experimental phases. In the first phase, we apply the emulation-based representation method described in Section 6.1 to learn representations for RNNs from Formal Languages and the Sequential MNIST dataset. The second phase is dedicated to predicting properties of the RNNs. These predictions are either based on representations learned in the first phase or derived from fully supervised models trained from scratch. We conduct the main experiments using 15 different random seeds for each model. The outcomes are presented as bootstrapped means with 95% confidence intervals. For RNNs trained on the Formal Languages dataset, performance is measured by the proportion of correctly generated strings (i.e., strings belonging to the language on which it was trained). For the Sequential MNIST dataset, we assess performance using standard digit classification validation accuracy. For the Flattened Weights and the Parameter Transformer encoders, the training data is augmented by randomly permuting the neurons of the input RNN.
@@ 6.1 Representation Learning
Layer-Wise Statistics 0.051 $\scriptstyle{(0.050,0.053)}$ 0.039 $\scriptstyle{(0.038,0.039)}$
@@
Figure 5: $f_{\theta}$’s original performance on formal language generation vs. the performance of $A_{\xi}$’s emulation based on $E_{\phi}(\theta)$ (validation data).
The six different encoder architectures are trained according to Objective 1. The hyperparameters of these encoders are selected to ensure a comparable number of parameters across all models. Each encoder generates a 16-dimensional representation $z$. An LSTM with two layers functions as the emulator $A_{\xi}$. The conditioning of $A_{\xi}$ on an RNN $f_{\theta}$ is implemented by incorporating a linear projection of the corresponding representation $z$ to the BOS token of the input sequence of $A_{\xi}$. More details and hyperparameters can be found in Appendix D.
Table 2 displays the emulation losses for the different encoder types on both the Formal Languages and the Sequential MNIST validation datasets. For Formal Languages, the interactive probing method outperforms the others. In the case of Sequential MNIST, the performance of both probing encoders is quite similar and superior to that of the mechanistic encoders. Figure 5 shows the emulation effectiveness of various RNNs from the Formal Languages dataset. We compute 16 equally spaced target performance values between the best and the worst performances in the datasets. For each target performance, we select the 15 RNNs with performances closest to each target. The x-position of each point represents the mean of the original performances, and the y-positions represent the mean of the emulated performances (the shaded areas give the 95% confidence intervals). The variance of the original performances for each point is relatively low, as can be seen from the shaded area around the identity line. For this dataset, only the interactive probing encoder yields representations that enable $A_{\xi}$ to effectively emulate the original RNN. Figure 11 in the Appendix shows the analogous results also for the Formal Languages training and test set, Figure 12 for the Sequential MNIST datasets.
Figure 7 (top) examines the structure of the embedding space created by the interactive probing encoder for the Formal Languages validation set. It visually demonstrates that our method successfully learns coherent representation spaces of RNN weights. Each point represents an RNN, reduced from $Z$ to two dimensions using principal component analysis. Each language forms its own cluster, with each cluster exhibiting a gradient representing the generation accuracy of the RNNs. We highlight one cluster, corresponding to the language $L_{-2,2,-2}$. In contrast, t-SNE dimensionality reduction (bottom) fails to represent such structures in RNN weights. In t-SNE reductions, the weights of the RNN models for a single language are scattered across the entire space. Similarly, Figure 7 shows the embedding space for the Sequential MNIST dataset, where the task involves rotating MNIST digits. This rotation, a continuous scalar, is evident in the embedding space, alongside the classification accuracy of each model. In the t-SNE plot, although small clusters are noticeable for the nine snapshots of each run, the overall embedding space lacks a coherent structure. Visualizations of embedding spaces for all encoder architectures and OOD test sets are presented in Appendix E.4.